OpenStudy (anonymous):

Solve "cos 3x = -1" algebraically given the domain [0,2pi)

OpenStudy (anonymous):

\[0\le x <2\pi\]

OpenStudy (campbell_st):

so 3x = cos^(-1) (-1) then 3x = pi so x = pi/3 cos is negative in 2nd and 3rd quadrants 2nd is pi - pi/3 3rd is pi + pi/3 then \[x = 4\pi/3, 2\pi/3\]